40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 19 23:44:00 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdfIMC_OLTF_20180320.pdf
    Reply  Tue Mar 20 16:53:03 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdf
       Reply  Tue Mar 20 21:08:03 2018, gautam, Update, IOO, IMC loop checkup 
          Reply  Tue Mar 20 22:44:45 2018, gautam, Update, IOO, IMC loop checkup 
             Reply  Wed Mar 21 15:52:45 2018, gautam, Update, IOO, MC error point calibration IMC_PDHdisc_20180321.pdfIMC_PDHdisc785_20180321.pdfIMC_oscope.pdf
                Reply  Wed Mar 21 17:31:21 2018, gautam, Update, IOO, MC error point calibration IMC_RF_noise_calib.pdf
                   Reply  Wed Mar 21 21:13:44 2018, gautam, Update, IOO, IMC noise budget IMC_RF_noise_calib.pdf
                      Reply  Tue Mar 27 01:39:44 2018, gautam, Update, IOO, PSL noise eater was off IR_ALS_20180326.pdf
                         Reply  Tue Mar 27 19:32:03 2018, arijit, Update, IOO, PSL noise eater was off MCREFL_TF.pdfMCREFL_SPECTRUM.pdf
                            Reply  Tue Mar 27 23:37:35 2018, gautam, Update, IOO, MC REFL PD removed IMG_6955.JPG
                               Reply  Wed Mar 28 21:31:39 2018, gautam, Update, IOO, MC REFL PD removed 
                            Reply  Wed Mar 28 21:47:37 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement LISO_fit_with_info.pdfLISO_fit.zip
                               Reply  Thu Mar 29 17:57:36 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement NOISE_EATER_On_OFF.pdf
                                  Reply  Fri Mar 30 06:14:31 2018, rana, Update, IOO, MCREFL_PD Optical response measurement 
                                     Reply  Fri Mar 30 22:37:36 2018, Kevin, Update, IOO, MCREFL_PD Optical response measurement REFLPD_DemodBoard.pdf
                      Reply  Thu Apr 5 04:36:56 2018, Kevin, Update, IOO, Coil driver noise MC3.pdfMC2.pdfCoilDriver.pdf
                         Reply  Thu Apr 5 12:13:18 2018, Koji, Update, IOO, Coil driver noise 
                            Reply  Fri Apr 6 21:39:09 2018, gautam, Update, IOO, MC2 suspension health checkup MC2_Freeswinging.pdf
                               Reply  Mon Apr 9 18:46:03 2018, gautam, Update, IOO, Further debugging MC_Freeswinging.pdfIMC_NB_20180409.pdf
                                  Reply  Tue Apr 10 14:28:44 2018, gautam, Update, IOO, Further debugging 
                                     Reply  Wed Apr 11 01:34:31 2018, gautam, Update, IOO, Activities today 
                                        Reply  Thu Apr 12 18:12:49 2018, gautam, Update, ALS, NPRO channels hijacked IR_ALS_20180412.pdfBeatMouthDrift.pngETMX_20180416.jpg
                                           Reply  Fri Apr 13 11:02:41 2018, gautam, Update, ALS, EX fiber polarization drift EX_pol_drift.png
                                              Reply  Fri Apr 13 17:56:26 2018, gautam, Update, ALS, Fibers switched out 
                                                 Reply  Sat Apr 14 14:42:09 2018, gautam, Update, ALS, Fibers switched out PSLdrift.png
                                                    Reply  Tue Apr 17 14:08:29 2018, gautam, Update, ALS, Fibers switched out PSL_fluctuations.png
                                                       Reply  Fri Apr 20 00:26:34 2018, gautam, Update, ALS, Fibers switched out 
                                                          Reply  Sat Apr 21 20:25:12 2018, gautam, Update, ALS, PSL fiber pickoff status PSL-beatMouthPickoff.png
                                  Reply  Thu Apr 19 00:03:51 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180409.pdfIMC_coils_20180418.pdf
                                     Reply  Thu Apr 19 17:15:35 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180419.pdfIMC_coilDriverNoises_20180419.pdf
                                        Reply  Sun May 20 23:43:50 2018, gautam, Update, IOO, Coil driver noise re-measurement MC1_coilDriver.pdfMC2_coilDriver.pdfMC3_coilDriver.pdfMC_coilDriverNoises.tgzActuationChainNoiseMeas.pdf
                                           Reply  Tue May 22 17:26:25 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                              Reply  Tue May 22 23:28:01 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                                 Reply  Wed May 23 17:58:48 2018, gautam, Update, IOO, MC1 Coil Driver pulled out MC1_monitorTF.pngMC1_ULnoise.pdf
                                                    Reply  Wed May 30 10:17:46 2018, gautam, Update, IOO, MC1 Coil Driver pulled out EM_pickup.pdfcoilDriverNoiseComparison.pdf
                            Reply  Fri Apr 6 22:23:53 2018, Kevin, Update, IOO, Coil driver noise MC2_coil_driver.pdf
Message ID: 13689     Entry time: Mon Mar 19 23:44:00 2018     Reply to this: 13690
Author: gautam 
Type: Update 
Category: IOO 
Subject: IMC loop checkup 

[koji, gautam]

  1. I began my investigations by measuring the voltage noise of the demod board outputs with an SR560 (G=100) and SR785 in the audio band.
    • Measurement made with PSL shutter closed, LO input of demod board driven with the nominal level of ~2.5dBm, RF input terminated.
    • Motivation was to look for any noise features.
    • Expected noise level is ~2nV/rtHz (Johnson noise of 50ohm) since there are no preamp electronics post SCLF-5 LP filter on this board.
    • Attachment #1 shows the results of the measurement for a few scenarios. Spectra only shown for the I channel but the Q channel was similar. The LO=+5dBm curve corresponds to driving the input at 5dBm with a marconi, to see if the label of the nominal level being +5dBm had anything to it.
    • The arches above 1kHz seemed suspicious to me, so I decided to investigate further.
  2. Looking at the IMC Demod board schematic, I I saw that there were 2 ERA-5SMs in there which are responsible for amplifying the 29.5MHz signal which serves as the LO to the oscillators.
  3. I pulled the demod board out and tested it on the electronics workbench. Koji and I couldn't make sense of the numbers we were seeing (all measurements made with Agilent analyzer and active FET probe with 100:1 attentuator).
  4. We eventually concluded that the ERA-5SMs were not exhibiting the expected gain of ~20dB. So we decided to swap these out.
  5. This sort of measurement is not ironclad as the output of the ERA-5SM goes to the mixer whose input impedance is dynamically varying as the diodes are switching. So even after replacing the suspect amplifiers with new ones, we couldn't make the numbers jive.
  6. We suspected that the new amplifiers were getting saturated. The 3dB saturation point for the ERA-5SM is spec'ed as ~19dBm.
    • We "measured" this by varying the input signal level and looking for deviation from linearity.
    • We saw that there was ~1dB compression for ~13dBm output from the ERA-5SM (after correcting for all attenuators etc). But this number may not be accurate in the absolute sense because of the unknown input impedance of the mixer.
    • Moreover, looking at the spec sheet for the mixer, JMS-1H, we found that while it wasn't ideal to operate the mixer with the LO level a few dBm below the expected +17dBm, it probably wasn't a show stopper.
  7. So we figured that we need 10dB of attenuation between the "nominal" LO input level of 2.7dBm and the input of the demod board in order to keep the ERA-5SM in the linear range. This has now been implemented in the form of an SMA attenuator.
  8. IMC locked straight away. But I noticed that PC drive RMS level was unusually large.
  9. I found that by increasing the "IN1" gain of the CM board to 12dB (from 2dB) and the "VCO gain" to 10dB (from 7dB), I could recover a transfer function with UGF ~140kHz and PM ~30degrees (need more systematic and wider span measurements of this, and also probably need to optimize the crossover gains). See Attachment #2 for my quick measurement tonight.
  10. Updated mcup to reflect these new gains. Tested autolocker a few times, seems to work okay.
  11. While it presumably was a good thing to replace the faulty amplifiers and prevent them from saturating, this work has not solved the primary problem of excess frequency noise on the PSL.

It is not clear to me why installing an attenuator to prevent amplifier saturation has necessitated a 10dB increase in the IN1 gain and 3dB increase in the VCO gain. Initially, I was trying to compensate for the gain by increasing the FSS "Common Gain" but in that setting, I found an OLTF measurement impossible. The moment I enabled the excitation input to the CM board, the lock was blown, even with excitation amplitudes as small as -60dBm (from the Agilent network analyzer).

This may also be a good opportunity to test out one of the aLIGO style FET mixer demod boards (recall we have 2 spare from the 4 that were inside the ALS demod box). I'm going to ask Steve to package these into a 1U chassis so that I can try that setup out sometime. From a noise point of view, the aLIGO boards have the advantage of having a x100 preamp stage straight after the mixer+LPF. We may need to replace the lowpass filter though, I'm not sure if the one installed is 1.9MHz or 5MHz.

I've left an SR785 and AG4395 near 1X2 in anticipation of continuing this work tomorrow.

Unrelated to this work - seems like the WFS DC and RF offsets had not been set in a while so I reset these yesterday. The frequent model restarts in recent times may mean that we have to reset these to avoid using dated offset values.

Attachment 1: IMC_RF_noise.pdf  326 kB  Uploaded Tue Mar 20 00:49:28 2018  | Hide | Hide all
IMC_RF_noise.pdf
Attachment 2: IMC_OLTF_20180320.pdf  150 kB  Uploaded Tue Mar 20 01:09:55 2018  | Hide | Hide all
IMC_OLTF_20180320.pdf
ELOG V3.1.3-